OpenStudy (anonymous):

Express the repeating decimal .99999... as a fraction in lowest terms

OpenStudy (anonymous):

1

OpenStudy (anonymous):

should be

OpenStudy (xapproachesinfinity):

Here is a useful technique that you can use in recurring decimal number let x=0.99999..... then 10x=9.9999... next subtract the second from the first that is 10x-x=9.9999....-0.9999.... you get 9x=9===>x=1 however you need to recognize the meaning of this! does 0.9999 actually equal 1?

OpenStudy (xapproachesinfinity):

from the very beginning you can say 0.9999 is so close to 1

OpenStudy (xapproachesinfinity):

it is a mind bugling hehehe

OpenStudy (xapproachesinfinity):

there are several proves that show 0.999....=1

OpenStudy (xapproachesinfinity):

here is something start with 1/9=0.1111..... then multiply both side by 9 so 9*(1/9)=9*0.1111.... 1=0.9999